OpenStudy (anonymous):

can somebody help me with a few GEOMETRIC problems, I really need help, PLEASE help me!!!!!!

OpenStudy (raden):

what ur question ?

OpenStudy (anonymous):

lol, Im trying to figure out how to put my question up here so people can answer it :)

OpenStudy (raden):

just type ur question in part "Ask Question" <------------------- if u have diagram, u can use utility "Draw" under

OpenStudy (anonymous):

ok hold on, I would draw it but it would take to long, but im still going to ask the question

OpenStudy (raden):

yup, try it :)

OpenStudy (anonymous):

ok

OpenStudy (raden):

dont forget to click "POST" :D

OpenStudy (anonymous):

If Z is the midpoint of RT what are x, RZ, and RT? A. x = 18, RZ = 134, and RT = 268 B. x = 22, RZ = 150, and RT = 300 C. x = 20, RZ = 150, and RT = 300 D. x = 20, RZ = 300, and RT = 150

OpenStudy (raden):

diagram ?

OpenStudy (anonymous):

can I just send you a link

OpenStudy (raden):

yes, u can

OpenStudy (raden):

well, we have to find the value of x, first RZ = ZT 8x-10 = 150 solve for x

OpenStudy (anonymous):

how did you get 8x-10 =150

OpenStudy (raden):

because Z is the midpoint, it means the length of RZ and ZT are same

OpenStudy (anonymous):

ok, so whats next

OpenStudy (raden):

just solve algebraic expression of 8x-10 = 150 what is x ?

OpenStudy (anonymous):

ok hold on

OpenStudy (raden):

have u tried ?

OpenStudy (anonymous):

im trying right now

OpenStudy (anonymous):

x is 20

OpenStudy (raden):

ok,that's right

OpenStudy (anonymous):

lets go!!!

OpenStudy (raden):

then find the length segement of RZ,